### v2

# An accurate vector optically pumped magnetometer with microwave-driven Rabi frequency measurements

An accurate vector optically pumped magnetometer with microwave-driven Rabi frequency measurements
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
Robust calibration of vector optically pumped magnetometers (OPMs) is a nontrivial task, but increasingly important for applications requiring high-accuracy such as magnetic navigation, geophysics research, and space exploration. Here, we showcase a vector OPM that utilizes Rabi oscillations driven between the hyperfine manifolds of 87Rb to measure the direction of a DC magnetic field against the polarization ellipse structure of a microwave field. By relying solely on atomic measurements—free-induction decay (FID) signals and Rabi measurements across multiple atomic transitions—this sensor can detect drift in the microwave vector reference and compensate for systematic shifts caused by off-resonant driving, nonlinear Zeeman (NLZ) effects, and buffer gas collisions. To facilitate dead-zone-free operation, we also introduce a novel Rabi measurement that utilizes dressed-state resonances that appear during simultaneous Larmor precession and Rabi driving (SPaR). These measurements, performed within a microfabricated vapor cell platform, achieve an average vector accuracy of 0.46 mrad and vector sensitivities down to 11 $\mu$rad$/\sqrt{\text{Hz}}$ for geomagnetic field strengths near 50 $\mu$T. This performance surpasses the challenging 1-degree (17 mrad) accuracy threshold of several contemporary OPM methods utilizing atomic vapors with an electromagnetic vector reference.
1 Introduction
Optically pumped magnetometers (OPMs) based on hot atomic vapors are increasingly important for both fundamental science and technological applications. These sensors utilize atomic spin precession, achieving record precision down to attotesla scales [1, 2, 3, 4]. Because the precession frequency is only sensitive to the magnetic field strength with respect to well-known atomic constants, OPMs are highly accurate and less dependent on the sensor orientation compared to directional magnetometers, such as fluxgates or superconducting quantum interference devices (SQUIDs), which measure individual magnetic field components. Consequently, OPMs are employed in a wide range of applications, including biosensing [5, 6, 7], contraband testing [8], and magnetic communications [9, 10, 11]. They also aid in dark matter searches [12, 13], probe fundamental symmetries beyond the energy scales reachable with modern particle colliders [14, 15], and serve as promising platforms for quantum-enhanced measurements that utilize squeezed light [16, 17] and macroscopic entanglement [18, 19].
Accurate vector magnetometry, however, remains a challenge for OPMs due to their inherent scalar operation, as they require an external reference, such as a coil system, to extract directional information. While scalar measurements are often sufficient, many applications benefit from vector measurements of the magnetic field including magnetic anomaly detection [20, 21], navigation [22, 23], space exploration [24, 25, 26], and geophysics [27]. For many of these applications accuracy is as crucial as precision. For instance, in planetary magnetosphere studies [24], enhanced vector accuracy would significantly improve the modeling of internal fields during satellite flybys. Because OPMs require external references to map magnetic field components, vector accuracy is often limited by machining tolerances, drift, and errors in modeling atomic signals.
To date the most common and accurate vector OPMs rely on scalar detection with respect to the magnetic fields produced by a coil system. The vector accuracy of these techniques are generally limited by the calibration accuracy of coil system parameters such as coil factors and coil pair orthogonality. These approaches include a directional varying reference field [28, 29], fast rotating fields [30], and low-frequency coil modulations [31, 32]. In particular, coil modulation is a well-established approach that has been implemented with a 4He OPM in the European Space Agency SWARM mission [33]. That sensor reached $10$ $\mu$rad directional accuracy after a calibration involving multiple sensor rotations [31].
Yet, vector magnetometry via scalar detection combined with a coil system presents certain technical challenges. The first challenge is that the vector sensitivity degrades with increased background magnetic field strength for the same modulation depth. Moreover, in many cases weak modulation fields are necessary due to power requirements, to prevent coupling to external objects, and slew-rate limitations of coil feedback electronics [34]. For instance, by employing modulation depths of around 18 $\mu$T, vector component sensitivities down to 0.4 pT$/\sqrt{\text{Hz}}$ have been achieved, as reported in [30]. Conversely, in the SWARM mission, vector component sensitivities are restricted to 1 nT$/\sqrt{\text{Hz}}$, a limitation attributed to the smaller modulation depths of 50 nT, as detailed in [33]. Another drawback is that coil modulation techniques offer limited insight from atomic measurements for monitoring environmental drifts in the vector reference (i.e. the coil system). As a result, achieving high accuracy can entail performing iterative calibrations without prior knowledge of whether drift occurred. This process often requires rotations of either the sensor or a large bias field, which can lead to intervals of unnecessary sensor downtime. Finally, a coil system enclosure around the OPM vapor cell increases the physical complexity of the apparatus and limits the overall capabilities for miniaturization [35].
In contrast, using an electromagnetic field as a vector reference, by coupling the atomic ensemble to its polarization structure or propagation direction, offers distinct advantages over coil modulation approaches. In particular, all-optical methods, such as electromagnetically induced transparency (EIT) [36, 37], nonlinear magneto-optical rotation (NMOR) [38, 39], and methods that detect spin projections on multiple laser beams [35, 40, 41, 42, 43], are attractive because they enable remote detection, are magnetically quiet, and are compatible with sensor miniaturization. Additionally, radio-frequency fields can be utilized in techniques such as double resonance atomic alignment [44, 45] and the Voigt effect [46]. Unlike coil system modulation techniques, these approaches do not gain vector information from a modulation in the magnetic field strength, with the exception of [40], but resonantly probe atomic transitions. Thus, by tuning the electromagnetic frequency their vector sensitivity does not degrade at large DC magnetic fields.
Despite these advantages, achieving vector accuracy better than 1-degree (17 mrad) with an electromagnetic reference is a nontrivial task due to challenges in modeling the nonlinear directional dependence of the atomic measurements [45, 42, 39, 47, 43]. Furthermore, most reports on vector OPMs using an electromagnetic reference focus solely on sensitivity characterization, leaving vector accuracy underexplored. Major contributing factors to modeling inaccuracy are sensitivity to complex decoherence effects from atomic collisions, imperfect optical pumping, and characterization errors of the electromagnetic reference in terms of spatial inhomogeneity, polarization structure, and time-dependent drifts.
Figure 1: Vector magnetometry using Rabi oscillation frequencies referenced to multiple microwave polarization ellipses (MPEs). (a) Energy-level diagram showing the four hyperfine transitions utilized for Rabi detection. Off-resonant microwave driving causes energy-level perturbations ($\Delta^{\mu\text{w}}_{\text{shift}}$) and buffer gas collisions cause a frequency shift in the hyperfine splitting ($\nu_{\text{bg}}$). (b) Generalized Rabi frequencies $\tilde{\Omega}_{m}^{m^{\prime}}$ are quantified by the difference between dressed-state energies $\lambda_{j}$ and $\lambda_{i}$. Off-resonant driving (red) from $\mathcal{B}^{(\alpha,\beta)}_{\pi}$, such as on the $\sigma^{+}$ transition, causes shifts in these dressed states. (c) The measured $\mathcal{B}_{\sigma^{+}}^{(\alpha,\beta)}$ spherical microwave component discretely sampled at different magnetic field directions for each of the three MPEs. (d) Measured Rabi-Chevron pattern. (e) The magnetometer apparatus.
In this article, we demonstrate sub-mrad vector magnetometry accuracy through Rabi frequency measurements driven by a microwave field, serving as an electromagnetic vector reference, across multiple atomic transitions. This is achieved through key information in the Rabi measurements to correct for collisional shifts and systematic errors from off-resonant driving, as well as facilitating robust calibration of the microwave field and detection of time-dependent drifts. In this technique, the Rabi measurements derive vector information from the polarization structure of the electromagnetic drive relative to a quantization axis defined by the unknown DC magnetic field. This approach could be extended to a wide range of electromagnetic drives and atomic systems. For example, extension to optical fields could be enabled through two-photon Raman transitions, or to RF fields by coupling magnetic sublevels within a hyperfine manifold. From a converse perspective, this work also highlights the ability of atomic systems to accurately map microwave field structures, with potential application in microwave circuit characterization [48, 49, 50], medical imaging [51, 52], and radar [53].
Previous experiments using Rabi frequency measurements to map DC electric field components [54] and magnetic field components [55] have only been implemented with ultracold atoms. These experiments showed limited sensitivity and accuracy arising from using small atomic ensembles, employing a single microwave drive, and neglecting effects from off-resonant driving  [54, 55]. The potential for Rabi measurements to compete with the precision of standard vector magnetometry techniques in vapor-cell platforms is evidenced by the coherence times of Rabi oscillations in vapor cells being similar to those of conventional free-induction decay (FID) measurements [56, 57]. In addition, hyperfine transitions exhibit larger magnetic couplings (e.g., $\mu_{m=1}^{m^{\prime}=2}/h\approx 17$ Hz/nT) than the gyromagnetic ratios $\gamma\approx 7$ Hz/nT characterizing Zeeman transitions.
For our implementation, we utilize microwave-driven Rabi oscillations across four hyperfine transitions of 87Rb [Fig 1(a,b)] to calibrate and reference the microwave polarization ellipse (MPE) structure of the driving field to an unknown magnetic field direction $(\alpha,\beta)$. For a given MPE structure ($\vec{\mathcal{B}}^{\mu\text{w}}$), the atomic quantization axis, defined by $(\alpha,\beta)$, gives rise to spherical microwave components $(\mathcal{B}_{\sigma^{+}}^{(\alpha,\beta)},\mathcal{B}_{\pi}^{(\alpha,\beta)},\mathcal{B}_{\sigma^{-}}^{(\alpha,\beta)})$ that are linked to the Rabi rates $\Omega_{m}^{m^{\prime}}$ coupling hyperfine states |$\ket{1,m}$ and |$\ket{2,m^{\prime}}$ through
$\Omega_{m}^{m^{\prime}}=\mu_{m}^{m^{\prime}}\mathcal{B}_{k}^{(\alpha,\beta)}/h,$ (1)
where $h$ is Planck’s constant, $\mu_{m}^{m^{\prime}}$ is the transition magnetic dipole moment, and $k=\sigma^{\pm},\pi$ denotes hyperfine transitions characterized by $m^{\prime}=m\pm 1$ and $m^{\prime}=m$ respectively. We indirectly detect the Rabi rates, and hence the spherical microwave components, through measurements of the generalized Rabi frequency $(\tilde{\Omega}_{m}^{m^{\prime}})$ that are well-approximated in the two-level formalism, apart from multi-level effects [Fig. 1(a,b)], as
$\tilde{\Omega}_{m}^{m^{\prime}}\approx\sqrt{(\Omega_{m}^{m^{\prime}})^{2}+(\Delta_{m}^{m^{\prime}})^{2}},$ (2)
where $\Delta_{m}^{m^{\prime}}=\nu_{\mu\text{w}}-\nu_{m}^{m^{\prime}}$ is the microwave detuning, $\nu_{\mu\text{w}}$ is the microwave frequency, and $\nu_{m}^{m^{\prime}}$ is the hyperfine transition resonance.
Initially, the MPE structure undergoes calibration using generalized Rabi frequencies $(\tilde{\Omega}_{m}^{m^{\prime}})$ measured over many known DC magnetic field orientations. Subsequently, these measurements, in conjunction with the directional maps [Fig. 1(c)] modeled with calibrated MPE parameters, determine the orientation of any magnetic field direction. The use of three distinct MPEs helps mitigate regions where the spherical microwave components, connected to Rabi rates through Eq. (1), weakly depend on the magnetic field direction (e.g. $\partial\mathcal{B}_{k}^{(\alpha,\beta)}/\partial\beta\approx 0$) as indicated in Fig. 1(c). Utilizing multiple MPE structures that do not all mutually have these regions for the same $(\alpha,\beta)$ enables sensitive vector measurements for all DC field directions.
We account for various systematic shifts in the Rabi oscillation measurements [Fig. 1(d)] that affect the accuracy of the directional MPE maps in Fig. 1(c). These shifts include the buffer-gas pressure shift ($\nu_{\text{bg}}$), nonlinear Zeeman shifts ($\Delta_{\text{NLZ}}$), and frequency shifts arising from off-resonant driving ($\Delta^{\mu\text{w}}_{\text{shift}}$) [Fig. 1(b)]. To correct for these effects, we model them using a full atom-microwave Hamiltonian to fit the Rabi data across multiple atomic transitions. Further, to mitigate systematic errors from microwave drift, we employ running MPE calibrations to track microwave field drifts over a 37.5-minute period, which are consistent with a drift observable derived from the Rabi measurements. We envision that such drift observables, discussed in Sec. 8, could be useful to avoid blind recalibration employed in high-accuracy vector magnetometers [58, 33].
Notably, these Rabi measurements use the same vapor cell parameters, such as temperature and buffer gas pressure, as those used in sensitive scalar OPMs. Thus, while the Rabi measurements determine the direction of a DC magnetic field, this approach can be combined with any scalar OPM measurement, such as free-induction decay (FID), to obtain the full magnetic field vector. Even so, we recently showed that Rabi measurements themselves can measure geomagnetic field strengths with improved accuracy over standard OPM measurements from 5 nT down to $0.6$ nT [59].
We also explore a novel scheme that extracts Rabi rates from resonances during simultaneous Larmor precession and Rabi driving (SPaR). While not as sensitive as Rabi oscillation measurements, SPaR obtains the optimal signal-to-noise ratio (SNR) for DC magnetic field directions where the Rabi oscillation SNR approaches zero. For the Rabi oscillation measurements, this probing deadzone occurs when the magnetic field is perpendicular to the probe beam. The SPaR approach leverages the fact that the Larmor precession signal is maximized at this Rabi probing deadzone. Notably, the existence of a deadzone is a feature to most OPM configurations using a single optical axis [60].
For geomagnetic fields near 50 $\mu$T we achieve 0.54 $\text{nT}/\sqrt{\text{Hz}}$ (11 $\mu$rad/$\sqrt{\text{Hz}}$) vector component (directional) sensitivity and 23 nT (0.46 mrad) vector component (directional) accuracy with Rabi oscillation measurements away from the Rabi probing deadzone. Using SPaR within the Rabi probing deadzone we achieve 7.3 $\text{nT}/\sqrt{\text{Hz}}$ (150 $\mu$rad/$\sqrt{\text{Hz}}$) sensitivity and 210 nT (4.3 mrad) accuracy. Although there is still room for engineering improvements to optimize microwave stability and reduce measurement deadtime, these results surpass the directional accuracy of many vector OPM techniques using an electromagnetic reference and demonstrate vector component sensitivities comparable to those of high-accuracy OPMs, such as the 4He OPM used in the SWARM mission [33].
2 Apparatus
[…]
Along the cell optical axis, parallel to $\hat{z}$, propagates a 795-nm elliptically polarized pump beam, tuned within a few gigahertz of the $D_{1}$ line, and a 1-mW probe beam blue-detuned by 170 GHz from the 780-nm $D_{2}$ line. The pump frequency and polarization are tuned to depopulate the $F=2$ manifold and enable strong Rabi signals across all hyperfine transitions, while still causing spin polarization for the FID measurement. Complete depopulation of the $F=2$ manifold is limited by the 5.6-GHz optical broadening from Rb-N2 collisions. A polarimeter detects the Faraday rotation $\theta_{F}=g\langle S_{z}\rangle+\theta_{0}$ of the probe beam expressed in terms of the macroscopic z component of the electron spin, a coupling coefficient $g$, and an offset $\theta_{0}$ [61]. In this scheme, atomic population dynamics affecting $S_{z}$ from the Rabi oscillations are nondestructively measured through $\theta_{F}$. Because the spin dynamics from microwave-driven Rabi oscillations occur along the DC magnetic field ($\vec{B}_{\text{dc}}$), this detection scheme exhibits a deadzone when $\vec{B}_{\text{dc}}$ is oriented perpendicular to the probe beam. The SPaR technique detailed in Sec. 6 enables Rabi rate measurements even in this probing deadzone.
The cell and microwave cavity are housed within a three-dimensional (3D) coil system that generates a programmable 50-$\mu$T magnetic field, defined by azimuthal and polar angles $\alpha$ and $\beta$. The coil system is further enclosed within three layers of mu-metal shielding to ensure magnetic isolation from the environment. Calibration of the coil system, detailed in Supplementary 1, defines the laboratory frame $\mathcal{L}=(x,y,z)$ and is essential for the MPE calibrations detailed in Sec. 5. Notably, outside of the MPE parameter calibration, with calibrated MPE parameters, vector magnetometry using Rabi measurements can be performed in the field independent of DC coil system drifts.
The DC magnetic field predicted from the coil system calibration also serves as an independent benchmark of the vector accuracy of the Rabi measurements. To account for coil system drift during the 37.5-minute vector data acquisition in our vector assessment in this work, we perform running calibrations of the coil system parameters post-processing from FID measurements described in Supplemental 1. From the FID calibration residuals and the consistency of calibrated coil system parameters across independent calibrations, the directional accuracy of the magnetic field programmed by the coil system, limited by FID heading errors [59] and nonlinearity in the coil current control, is estimated to be within 50 $\mu$rad. This establishes a lower limit on the accuracy with which vector magnetometry from Rabi measurements can be benchmarked. Further details behind the coil system design and calibration are discussed in Supplemental 1.
3 Measurement Protocol
The measurement protocol for conducting Rabi oscillation and SPaR measurements across the four hyperfine transitions [Fig. 1(a)] and three MPE configurations [Fig. 1(c)], along with FID measurements, is diagrammed in Fig. 2. For the Rabi oscillation measurements, we initialize the atomic ensemble with adiabatic optical pumping (AOP) to eliminate Larmor precession by aligning the macroscopic atomic spin along the magnetic field $\vec{B}_{\text{dc}}$ as described in [59]. As depicted in Fig. 2(a), this involves a 100-mW pump pulse lasting for 50 $\mu$s, followed by a linear ramp-off of the pump power over an additional 50 $\mu$s. In contrast, FID and SPaR measurements utilize a single 100-$\mu$s pulse of pump light at 400 mW to polarize the atomic spins along the pump beam. After pumping, the microwave field from one of the three MPEs is turned on for 0.85 ms and 1.85 ms for Rabi oscillation and SPaR measurements respectively. We drive both of the Rabi and SPaR signals for each hyperfine transition highlighted in Fig. 1(a) using a single microwave frequency $\overline{\nu}_{m}^{m^{\prime}}$ that is near-resonant to the unperturbed hyperfine transition frequency $\nu_{m}^{m^{\prime}}$. The specific microwave frequencies are $\overline{\nu}_{-1}^{-2}=6833.7201$ MHz, $\overline{\nu}_{-1}^{-1}=6834.0701$ MHz, $\overline{\nu}_{1}^{1}=6835.4701$ MHz, and $\overline{\nu}_{1}^{2}=6835.8204$ MHz.
Figure 2: (a) Rabi oscillation and SPaR pulse sequences measured at $(\alpha,\beta)=(0\degree,57\degree)$. (b) A timing diagram of the Rabi, SPaR, and FID measurement sequences utilized to benchmark vector accuracy and sensitivity. All Rabi, SPaR, and FID measurements are taken within a $\Delta t_{R}+\Delta t_{\text{FID}}=430$ ms period. Magnetic field rotations (blue arrow) and deactivation of electrical heating are initiated 1 s and 40 ms before each measurement period to minimize Eddy currents during acquisition. All 12 Rabi and SPaR measurements, corresponding to the three MPE configurations and four hyperfine transitions, are each repeated 10 times.
The Rabi, SPaR, and FID measurements are made over a range of DC magnetic field orientations [Fig. 2(b)]. The sequence of $\vec{B}_{\text{dc}}$ orientations consist of a random direction $(\alpha,\beta)$ interspersed with two repeated field directions used for monitoring microwave drifts. These measurements allow for the evaluation of vector sensitivity and accuracy across all magnetic field directions and enable ongoing MPE calibrations to correct for microwave field drift. At each $\vec{B}_{\text{dc}}$ direction, Rabi oscillation and SPaR measurements for each $\overline{\nu}_{m}^{m^{\prime}}$ and MPE configuration are repeated 10 times over a $\Delta t_{R}=400$ ms period followed by 10 repeated FID measurements over a $\Delta t_{\text{FID}}=30$ ms period [Fig. 2(b)]. The 400 ms Rabi measurement sequence is initiated twice before capturing the final Rabi-SPaR-FID sequence to allow time for the microwave components to thermally stabilize.
To prevent Eddy currents during measurements, the cavity heaters are deactivated 40 ms before the final sequence, and the magnetic field rotations are completed 1 second prior. The cooling that occurs due to the deactivation of the cavity heaters during the Rabi-SPaR-FID sequence necessitates a dead time after each sequence to allow the heaters to restore the cavity temperature to $100$$\degree$°C. This dead time restricts the repetition rate of the measurement sequences to $\Delta t=2.5$ seconds and also limits the speed at which MPE calibrations can be performed, as detailed in Sec. 5. Future systems that limit this deadtime will be even less sensitive to drifts during this calibration period.
4 Generalized Rabi Frequency Modeling
Generalized Rabi frequencies are influenced by various factors [Fig. 1(a,b)], including off-resonant driving, buffer gas collisions, and nonlinear Zeeman shifts, which can constrain the ultimate accuracy of the vector measurements. To account for these frequency shifts we model generalized Rabi frequencies in terms of dressed state energies $\lambda_{j}$ and $\lambda_{i}$ that correspond to the pair of states coupled by the microwave field [Fig. 1(b)]. This relationship is expressed as [59]
$\tilde{\Omega}_{m}^{m^{\prime}}=\delta\lambda_{m}^{m^{\prime}}\equiv(\lambda_{j}-\lambda_{i})/h.$ (3)
Specifically, $\lambda_{j}$ and $\lambda_{i}$ are eigenvalues of the atom-microwave Hamiltonian
$\begin{split}H=&\mathcal{M}\big{[}(A+h\frac{\nu_{\text{bg}}}{2})\mathbf{S}\cdot\mathbf{I}+\mu_{B}(g_{s}S_{z}+g_{i}I_{z})B\big{]}\mathcal{M}^{\dagger}-\\ &I_{2}h\nu_{\mu\text{w}}+\sum_{|m-m^{\prime}|\leq 1}\frac{h}{2}\Big{[}\ket{\overline{2,m^{\prime}}}\Omega_{m}^{m^{\prime}}\bra{\overline{1,m}}+\text{h.c.}\Big{]}\end{split}$ (4)
The Hamiltonian $H$ in Eq. (4) is defined within an atom-frame $\mathcal{A}=(x_{a},y_{a},z_{a})$ where the $z_{a}$-direction is aligned with the magnetic field direction $(\alpha,\beta)$ with respect to the lab frame $\mathcal{L}$ [Fig. 1(e)]. The first line describes the hyperfine and Zeeman interaction and is characterized by electronic $\mathbf{S}=(S_{x},S_{y},S_{z})$ and nuclear $\mathbf{I}=(I_{x},I_{y},I_{z})$ spin operators defined in the |$\ket{F,m_{F}}$ basis, the magnetic dipole hyperfine constant $A$, electron $g_{s}$ and nuclear $g_{i}$ Landé g-factors, the Bohr magneton $\mu_{B}$, the DC magnetic field strength $B=|\vec{B}_{\text{dc}}|$, and the buffer gas frequency shift $\nu_{\text{bg}}\approx 88$ kHz. The magnetic field strength $B\approx 50$ $\mu$T is known from FID measurements, and the pressure shift $\nu_{\text{bg}}$ is known from an independent characterization discussed in Supplementary 1. The second line describes the atom-microwave coupling where $I_{2}$ is the $F=2$ identity operator. We work in a modified hyperfine basis |$\ket{\overline{F,m}}=\mathcal{M}\ket{F,m}$ with $\mathcal{M}$ being defined as the operator that diagonalizes the hyperfine and Zeeman terms to preserve NLZ effects during the rotating-wave approximation (RWA).
The spherical microwave components that determine the Rabi rates $\Omega_{m}^{m^{\prime}}$ through Eq. 1 are defined within the atom-frame $\mathcal{A}$ as
$\mathcal{B}_{k}^{(\alpha,\beta)}=R_{y}(-\beta)R_{z}(-\alpha)\vec{\mathcal{B}}\cdot\epsilon_{k}$ (5)
where $\vec{\mathcal{B}}=(\mathcal{B}_{x}e^{-i\phi_{x}},\mathcal{B}_{y}e^{-i\phi_{y}},\mathcal{B}_{z})$ is a complex phasor written in terms of microwave amplitudes $(\mathcal{B}_{x},\mathcal{B}_{y},\mathcal{B}_{z})$ and relative phases $(\phi_{x},\phi_{y})$ defined in the lab frame $\mathcal{L}$. This phasor describes any microwave field through $\mathcal{B}(t)=\frac{1}{2}[\vec{\mathcal{B}}e^{-i\omega_{\mu\text{w}}t}+\vec{\mathcal{B}}^{\star}e^{i\omega_{\mu\text{w}}t}]$. Spherical projection operators $\epsilon_{\pm}=\{\frac{1}{\sqrt{2}},\mp\frac{i}{\sqrt{2}},0\}$ and $\epsilon_{\pi}=\{0,0,1\}$ are also defined within the atom-frame $\mathcal{A}$. The phasor $\vec{\mathcal{B}}$ is rotated into the atom-frame through 3D rotation operators $R_{y,z}$ defined with respect to the lab frame $\mathcal{L}$.
For clarity, throughout this article we denote the phasor $\vec{\mathcal{B}}$ and generalized Rabi frequency $\tilde{\Omega}_{m}^{m^{\prime}}$ corresponding to the $p^{\text{th}}$ MPE configuration, where $p=1,2,3$, as $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ and $(\tilde{\Omega}_{m}^{m^{\prime}})_{p}$ respectively. This notation also specifies the microwave frequency $\nu_{\mu\text{w}}=\overline{\nu}_{m}^{m^{\prime}}$ associated with driving a specific hyperfine transition.
5 MPE Calibration and Long-Term Drift
An essential step to performing vector magnetometry with the Rabi measurements is to calibrate the microwave field parameters. Because of the microwave frequency dependence of the cavity modes, 12 phasors $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ must be calibrated, corresponding to 60 microwave parameters. We fit each $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ using Eq. (3) over $N=12$ DC field directions by minimizing the following cost function
$(r_{m}^{m^{\prime}})_{p}=\sum_{j=1}^{N}(w_{m}^{m^{\prime}})_{p}\Big{(}\delta\lambda_{m}^{m^{\prime}}((\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p},\alpha,\beta)-(\tilde{\Omega}_{m}^{m^{\prime}})_{p}\Big{)}^{2}\Big{|}_{(\alpha_{j},\beta_{j})}.$ (6)
We reiterate that during these MPE calibrations and benchmarking, the angles $(\alpha,\beta)$ are determined with an accuracy of $50$ $\mu$rad from running coil system calibrations discussed in Sec. 2. In Eq. (6) we explicitly denote the magnetic field direction and microwave field dependence of $\delta\lambda_{m}^{m^{\prime}}$. The generalized Rabi frequencies $(\tilde{\Omega}_{m}^{m^{\prime}})_{p}$ and corresponding weights $(w_{m}^{m^{\prime}})_{p}=1/(\delta(\tilde{\Omega}_{m}^{m^{\prime}})_{p})^{2}$ are evaluated from the mean and variance $(\delta(\tilde{\Omega}_{m}^{m^{\prime}})_{p})^{2}$ of 10 repeated Rabi oscillation measurements.
Figure 3: (a) The sequence of $\vec{B}_{\text{dc}}$ directions $(\alpha,\beta)$ consisting of random directions (black circles) interspersed with repeated directions $(0\degree,57\degree)$ and $(135\degree,57\degree)$. While all $\vec{B}_{\text{dc}}$ directions in this sequence are utilized in the MPE calibrations, vector measurements are only evaluated at the random $(\alpha,\beta)$. Conversely, the repeated field directions (gray and white circles) are used to consistently monitor microwave drifts. To avoid overfitting, the microwave phasor $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ at each random $\vec{B}_{\text{dc}}$ direction in this sequence is determined from the weighted average of the two phasors $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p,\text{prev}}$ and $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p,\text{sub}}$ extracted from the MPE calibrations that consist of the $N=12$ directions immediately before and after the random $(\alpha,\beta)$ direction. The calibration length $N$ is limited by the finite time $\Delta t=2.5$ s that is dominated by measurement dead time reserved for cavity heating. (b) Measured MPEs (solid lines) and their projections onto the coordinate planes (dotted lines). The nT-scale variation of the microwave field components arising from the frequency dependence of the cavity modes are not discernible in this plot. (c) Residuals of the $\sigma^{+}$ generalized Rabi frequency $(\tilde{\Omega}_{m=1}^{m^{\prime}=2})_{p=2}$ at random DC magnetic field directions with and without MPE recalibration. (d) Drift of the $\mathcal{B}_{x}$ microwave component evaluated at different $\nu_{\mu\text{w}}=\overline{\nu}_{m}^{m^{\prime}}$. Here, MPE 2 measurements are used as a representative example in (c,d).
To account for MPE drifts, we make Rabi measurements over an $(\alpha,\beta)$ sequence consisting of random $\vec{B}_{\text{dc}}$ directions interspersed with two repeated constant directions [Fig. 3(a)]. At each random $(\alpha,\beta)$, we estimate $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ by using Rabi measurements taken at the 12 previous directions to calibrate a phasor $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p,\text{prev}}$ with squared residual error (Eq. (6)) $(r_{m}^{m^{\prime}})_{p,\text{prev}}$ and the subsequent 12 directions to calibrate $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p,\text{sub}}$ with the corresponding error $(r_{m}^{m^{\prime}})_{p,\text{sub}}$, and calculate the weighted average
$(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}=\frac{((r_{m}^{m^{\prime}})_{p,\text{sub}}(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p,\text{prev}}+(r_{m}^{m^{\prime}})_{p,\text{prev}}(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p,\text{sub}})}{((r_{m}^{m^{\prime}})_{p,\text{prev}}+(r_{m}^{m^{\prime}})_{p,\text{sub}})}.$ (7)
This approach avoids overfitting by ensuring that the calibrated MPE parameters in $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ are validated exclusively against Rabi measurements taken from magnetic field directions not used during the calibration process.
Over 37.5 minutes, we observe nT-scale drifts of the microwave field components and mrad-scale drifts in the phases as exemplified for MPE 2 in Fig. 3(d). The microwave field drift is dominated by the contraction of the microwave cavity modes due to a few °$\degree$°C cooling of the microwave cavity. Details on monitoring the temperature drift from the Rabi dephasing rate and accounting for temperature-dependent shifts in $\nu_{\text{bg}}$ are discussed in Supplementary 1. For these calibrations, we found that it was sufficient to recalibrate only the relative phase $\delta\phi=\phi_{y}-\phi_{x}$ by leaving $\phi_{x}$ fixed after an initial calibration instead of resolving for both $\phi_{x}$ and $\phi_{y}$. We note that most of this 37.5 minute duration is spent heating the cavity and not taking measurements. Calibrating all MPE parameters across 12 random $\vec{B}_{\text{dc}}$ directions currently requires 132 seconds. A promising approach to eliminate this duration in future work involves the use of continuous laser heating [62], which would not interfere with atomic measurements. By disregarding technical deadtime, the calibration time could be reduced to 1.8 seconds, as the total measurement duration for Rabi oscillation data is 150 ms. This duration could be further shortened to a few hundred milliseconds by reducing the MPE model parameters, for instance, by eliminating the frequency dependence of the cavity modes with improved microwave engineering. Shortening the calibration time would also help minimize systematic errors caused by potential drifts in the DC background magnetic field during MPE calibration.
To assess the calibration accuracy, we show residuals of $\sigma^{+}$ generalized Rabi frequencies $(\tilde{\Omega}_{m=1}^{m^{\prime}=2})_{p=2}$ against the $\delta\lambda_{m=1}^{m^{\prime}=2}$ model utilizing recalibrated MPE 2 parameters in Fig. 3(c) as a representative example. The residuals are only evaluated over random magnetic field orientations. Over 37.5 minutes these residuals exhibit a standard deviation of $\sigma=13$ Hz, which corresponds to $\sigma=0.8\text{ nT}$ fluctuations for the spherical microwave component $\mathcal{B}_{\sigma^{+}}^{(\alpha,\beta)}$. These fluctuations are approximately three times greater than the anticipated $\sigma=5$ Hz ($0.3$ nT), a figure expected due to the limited calibration length of $N=12$ (see Supplement 1). This discrepancy suggests that the systematic errors in the calibrated MPE parameters are estimated to be $\delta\mathcal{B}_{x}=0.4$ nT, $\delta\mathcal{B}_{y}=0.3$ nT, $\delta\mathcal{B}_{z}=0.2$ nT, and $\delta(\phi_{y}-\phi_{x})=0.3$ mrad. These calibration errors are consistent with the directional accuracies deduced in the vector magnetometry evaluation in Sec. 7.
6 Simultaneous Precession and Rabi (SPaR)
For the single optical-axis geometry utilized in this work, Rabi oscillations, and corresponding vector measurements, exhibit a deadzone when $\beta\rightarrow 90\degree$ [Fig. 4(a)]. In this section we discuss how SPaR is an effective method to detect Rabi rates $\Omega_{m}^{m^{\prime}}$ within this probing deadzone through precessional frequency components that appear around the Larmor frequency $\nu_{L}$ labeled $\nu_{\text{low}}$ and $\nu_{\text{hi}}$ in Fig. 4(a). Like FID measurements, the amplitude of these SPaR peaks are maximal when $\beta=90\degree$.
[…]
Figure 4: Rabi rate detection using SPaR. (a) Directional dependence of Rabi oscillation and SPaR signal amplitude. This plot shows the fast Fourier transform (FFT) of SPaR time-domain signals like the one shown in Fig. 2(a). (b) Energy-level diagram in the rotating frame near the $\sigma^{+}$ transition frequency. Coupling between microwave dressed states (green) with Zeeman sublevels (black) are indicated by double-sided arrows that indicate $\nu_{\text{hi}}$ (long, dashed) and $\nu_{\text{low}}$ (short, solid) frequency components. (c) Generalized Rabi frequency measurements using Rabi oscillations (green) and SPaR (black). SPaR measurements are shifted by the values in parenthesis. (d) Measured SPaR spectrum for the $\sigma^{+}$ transition. The top plot shows the real part of the SPaR FFT at the $\sigma^{+}$ transition frequency $\nu_{\mu\text{w}}=\overline{\nu}_{m=1}^{m^{\prime}=2}$, both with and without the broadened Larmor resonances occuring at $\nu_{\text{L}}$. A green dashed line represents a fit applied to the SPaR FFT after removing the Larmor resonances. The bottom plot shows the absolute value of the SPaR FFT over a range of microwave frequencies centered about $\nu_{\mu\text{w}}=\overline{\nu}_{m=1}^{m^{\prime}=2}$, with the broadened Larmor resonances retained. Vertical lines indicate $\nu_{\text{low}}$, $\nu_{\text{hi}}$, and $\nu_{\text{L}}$ precessional frequencies calculated from the eigenstates of $H$.
To fit Rabi rates $\Omega_{m}^{m^{\prime}}$ with SPaR, we begin by estimating the three spherical microwave components $\mathcal{B}_{k}^{(\alpha,\beta)}$, defined in Eq. (5), from the $\sigma^{+}$ and $\sigma^{-}$ SPaR spectra. We do this by first estimating the centers of the $\nu_{\text{low}}$ and $\nu_{\text{hi}}$ peaks in the SPaR FFT signal using a peak-finding algorithm. Then, we approximate $\nu_{\text{hi}}-\nu_{\text{low}}\approx\Omega_{m}^{m^{\prime}}$ along with Eq. (1) to estimate $\mathcal{B}_{\sigma^{\pm}}^{(\alpha,\beta)}$. From the total microwave field amplitude $|\vec{\mathcal{B}}|$, known from MPE calibrations described in Sec. 5, $\mathcal{B}_{\pi}^{(\alpha,\beta)}$ is calculated using $|\mathcal{B}_{\pi}^{(\alpha,\beta)}|^{2}=|\vec{\mathcal{B}}|^{2}-|\mathcal{B}_{\sigma^{+}}^{(\alpha,\beta)}|^{2}-|\mathcal{B}_{\sigma^{-}}^{(\alpha,\beta)}|^{2}$. By assuming that the spherical components are positive instead of complex numbers, we can estimate all Rabi rates in Eq. (4) with 1 kHz accuracy, and further, estimate all eigenvalues $\lambda_{j}$ of $H$. We linearize these eigenvalues about these estimates to enable faster computation during SPaR fitting that is described in Supplement 1.
Next, we remove the broadened Larmor resonance $\nu_{L}\approx 350$ kHz appearing in each SPaR spectra (Supplement 1) and use the following fitting equation to fit the $\sigma^{\pm}$ SPaR spectra
$\mathcal{R}e[\text{FFT}]=\sum_{j\in\{\text{low,hi}\}}a_{j}\frac{\text{cos}[\phi_{j}]-\text{sin}[\phi_{j}](f-\overline{\nu}_{j}+f_{\text{shift}})}{(f-\overline{\nu}_{j}+f_{\text{shift}})^{2}+w_{j}^{2}/4}$ (8)
where $\overline{\nu}_{j}$ is the mean frequency of the two $\nu_{j}^{(1)}$ and $\nu_{j}^{(2)}$ frequency components, $f_{\text{shift}}$ is a phenomenological frequency shift, $\phi_{j}$ are phase shifts, and the strength and broadening of this signal is given by amplitudes $a_{j}$ and linewidths $w_{j}\approx 2$ kHz. We use the mean frequency $\overline{\nu}_{j}$ instead of precisely modeling the lineshape of each of the $\nu_{j}^{(1)}$ and $\nu_{j}^{(2)}$ to avoid overfitting. During the $\sigma^{+}$ SPaR fitting, $\mathcal{B}_{\sigma^{+}}^{(\alpha,\beta)}$ is the only free parameter, where $\mathcal{B}_{\sigma^{-}}^{(\alpha,\beta)}$ and $\mathcal{B}_{\pi}^{(\alpha,\beta)}$ are held at their initial estimates. Correspondingly, $\mathcal{B}_{\sigma^{-}}^{(\alpha,\beta)}$ and $\mathcal{B}_{\pi}^{(\alpha,\beta)}$ are the only free parameters during SPaR fitting of the $\sigma^{-}$ and $\pi$ transitions respectively.
The accuracy of the SPaR Rabi fits is within a few hundred Hz [Fig. 4(c)] due to uncertainty in the relative amplitudes of the unresolved frequency components that make up $\nu_{\text{low}}$ and $\nu_{\text{hi}}$. Despite this, over a 17-minute period, we observe changes in Rabi rates using SPaR that align with those measured from Rabi oscillations as illustrated in Fig. 4(c). These observations occur at a magnetic field orientation $(\alpha,\beta)=(135\degree,57\degree)$, which is intermediary to the Rabi and SPaR deadzones.
7 Vector Magnetometry Evaluation
The evaluation of the accuracy and sensitivity of the vector measurements, conducted over 300 random $\vec{B}_{\text{dc}}$ orientations, are depicted in Fig. 5. In the case of Rabi oscillation data, we exclude field directions that have polar angles within $90\degree\pm 5\degree$, whereas for SPaR, the evaluation is specifically focused on polar angles that fall within $90\degree\pm 10\degree$. Despite the periodicity of $\mathcal{B}^{(\alpha,\beta)}_{k}$ with respect to DC magnetic field direction $(\alpha,\beta)$ [Fig. 1(c)], the Rabi measurements across three MPEs find unique solutions for $(\alpha,\beta)$. See Supplemental 1 for further discussion on the uniqueness of the field orientation deduced from the Rabi measurements.
For the Rabi oscillation data, we solve for the DC magnetic field direction $(\alpha,\beta)$ by fitting the generalized Rabi frequency $(\tilde{\Omega}_{m}^{m^{\prime}})_{s}$ to the eigenvalue model $\delta\lambda_{m}^{m^{\prime}}$ with the cost function
$r_{\text{rabi}}^{(\alpha,\beta)}=\sum_{p=1}^{3}\sum_{m,m^{\prime}}(\tilde{w}_{m}^{m^{\prime}})_{p}\Big{(}\delta\lambda_{m}^{m^{\prime}}((\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p},\alpha,\beta)-(\tilde{\Omega}_{m}^{m^{\prime}})_{p}\Big{)}^{2}.$ (9)
We note that the microwave parameters $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ are obtained from the calibration described in Sec. 5. For the SPaR data, we solve for $(\alpha,\beta)$ by fitting the measured Rabi rate $(\Omega_{m}^{m^{\prime}})_{p}$ to the spherical microwave component (Eq. (5)) with the cost function
$\begin{split}r_{\text{spar}}^{(\alpha,\beta)}&=\sum_{p=1}^{3}\sum_{m,m^{\prime}}(w_{m}^{m^{\prime}})_{p}\times\\ &\Big{(}\Big{|}R_{y}(-\beta)R_{z}(-\alpha)\frac{\mu_{m}^{m^{\prime}}(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}}{h}\cdot\epsilon_{k}\Big{|}-(\Omega_{m}^{m^{\prime}})_{p}\Big{)}^{2}.\end{split}$ (10)
The absolute value in Eq. (10) emphasizes the approximation, only during SPaR fitting, that the spherical microwave components and Rabi rates are positive numbers. For both, Eq. (9) and Eq. (10), the pairs of magnetic sublevels $m,m^{\prime}$ are only summed over the four hyperfine transitions in Fig. 1(a). The weights $(\tilde{w}_{m}^{m^{\prime}})_{p}=(\delta(\tilde{\Omega}_{m}^{m^{\prime}})_{p})^{-2}((r_{m}^{m^{\prime}})_{p})^{-1}$ in Eq. (9) are given by the fitting errors $\delta(\tilde{\Omega}_{m}^{m^{\prime}})_{p}$ of the generalized Rabi frequencies and further weighted by the MPE squared residual error $\overline{r}_{m}^{m^{\prime}}=((r_{m}^{m^{\prime}})_{p,1}+(r_{m}^{m^{\prime}})_{p,2})/2$, defined in terms of $(r_{m}^{m^{\prime}})_{p,1}$ and $(r_{m}^{m^{\prime}})_{p,2}$ introduced in Sec. 5. The weights $(w_{m}^{m^{\prime}})_{p}$ in Eq. (10) are similarly defined.
To avoid issues arising from $\alpha$ becoming undefined as $\beta\rightarrow 0$, accuracy and sensitivity evaluations are conducted by mathematically rotating the DC magnetic field $\vec{B}_{m}$ onto the z-axis of the lab frame $\mathcal{L}$ through
$(\delta B_{x},\delta B_{y},B_{z})=R_{y}(-\beta)R_{z}(-\alpha)\vec{B}_{m}.$ (11)
The rotation in Eq. (11) utilizes the field direction $(\alpha,\beta)$ predicted by a coil system calibration with 50 $\mu$rad accuracy (see Supplement 1). As mentioned in Sec. 2, running coil system calibrations are used throughout the vector measurements to eliminate errors due to coil system drifts. Up to the accuracy of the coil system calibration, residual transverse components $\delta B_{x}$ and $\delta B_{y}$ are due to vector inaccuracies. These transverse components are converted into a directional accuracy through
$\delta\theta=\text{tan}^{-1}\Big{(}\frac{1}{B}\sqrt{\delta B_{x}^{2}+\delta B_{y}^{2}}\Big{)},$ (12)
where $B=|\vec{B}_{m}|$. This directional accuracy is represented by the shaded circular arcs in Fig. 5(a,b).
Figure 5: Analysis of vector accuracy and sensitivity using either Rabi oscillation and SPaR measurements. (a) and (b) show the vector accuracy, and (c) and (d) show the vector sensitivity for Rabi and SPaR techniques evaluated over 300 and 80 random field directions respectively.
The mean directional accuracies of the Rabi oscillation and SPaR vector data are $0.46$ mrad ($0.026^{\circ}$) and $4.3$ mrad ($0.25^{\circ}$) respectively. These accuracies are consistent with the measured MPE calibration residuals from Rabi oscillation measurements at the 10 Hz scale [Fig. 3(c)] and the 100 Hz scale systematics in the SPaR measurements [Fig. 4(d)]. Rabi frequency systematic errors at the 10 Hz scale may partially be due to imperfect microwave drift compensation, but can also be due to a variety of sources stemming from microwave phase noise, DC magnetic field spatial gradients, and microwave spatial inhomogeneity discussed further in Supplement 1.
We characterize the vector component sensitivity in terms of the fluctuations of $\delta B_{x}$ and $\delta B_{x}$ in Eq. (11) through
$S_{B_{i}}=\sigma_{B_{i}}\sqrt{2t_{m}},$ (13)
where $i=x,y$, $\sigma_{B_{i}}$ is the standard error evaluated from 10 repeated measurements of the $\delta B_{i}$ components, and $t_{m}$ is the total measurement time. For the Rabi oscillation data, the total measurement duration is $t_{m}=150$ ms, while for SPaR measurements the duration is $t_{m}=270$ ms. Included in each $t_{m}$ is the FID measurement time lasting 30 ms. The directional sensitivity is evaluated in terms of the geometric mean of $S_{B_{x}}$ and $S_{B_{y}}$ through
$S_{\theta}=\text{tan}^{-1}\Big{(}\frac{1}{B}\sqrt{\delta S_{B_{x}}^{2}+S_{B_{y}}^{2}}\Big{)}.$ (14)
With this sensitivity metric, the Rabi oscillation vector measurements reach a mean directional sensitivity of $49$ $\mu$rad/$\sqrt{\text{Hz}}$ evaluated over random field directions, with the best sensitivity reaching $11$ $\mu$rad/$\sqrt{\text{Hz}}$. As examined in Supplemental 1, the measured directional sensitivity is about three times worse than expected if the noise in the Rabi measurements were purely due to polarimeter noise. Suboptimal performance below this detector limit is attributed to shot-to-shot amplitude fluctuations of the microwave field. A summary of the accuracy and sensitivity metrics for the Rabi oscillation and SPaR data are reported in Table 1.
As a sensitivity and bandwidth comparison, the vector component sensitivities reported in Table 1 are at the same 1 nT$/\sqrt{\text{Hz}}$ level reached in the Helium SWARM mission  [31]. With a measurement duration of $t_{m}=150$ ms, the vector magnetometer would achieve a bandwidth of 3.3 Hz if not for the currently implemented deadtime required for cavity heating and eddy current dissipation. For perspective, the vector mode of the ESA Swarm 4He magnetometer, operating in near-Earth orbit, has a bandwidth of 0.4 Hz [27]. While this is not a direct comparison due to a variety of apparatus differences such as the smaller cell volume employed in this work, it demonstrates comparable vector sensitivity and bandwidth to high accuracy OPM implementations. Future advancements could include implementing full 3D control of the microwave field, moving beyond planar MPEs. This would enable the strategic selection of MPEs to optimize vector sensitivity across all field directions. Additionally, as outlined in Sec. 5, continuous laser heating [62] offers a promising approach to eliminating deadtime associated with cavity heating. Replacing the microwave cavity with a dielectric resonator platform is another viable path to mitigate eddy current effects while reducing the sensor’s overall size and weight.
Table 1: A summary of the accuracy and sensitivity metrics for the Rabi oscillation and SPaR measurements. The first two rows detail the vector component and directional accuracy, while the last four rows describe vector component and directional sensitivity. Brackets delimited-⟨$\langle$..$\rangle$ … ⟩ indicate an average over random DC magnetic field directions, while min()(...)( … ) represents the minimum measured value.
Vector Metric Rabi SPaR
$\Big{\langle}\sqrt{\delta B_{x}^{2}+\delta B_{y}^{2}}\Big{\rangle}$ 23 nT 210 nT
$\big{\langle}\delta\theta\big{\rangle}$ 0.46 mrad 4.3 mrad
$\Big{\langle}\sqrt{\delta S_{B_{x}}^{2}+\delta S_{B_{y}}^{2}}\Big{\rangle}$ 2.5 nT/$\sqrt{\text{Hz}}$ 25 nT/$\sqrt{\text{Hz}}$
$\big{\langle}S_{\theta}\big{\rangle}$ 49 $\mu$rad/$\sqrt{\text{Hz}}$ 490 nT/$\sqrt{\text{Hz}}$
min$\Big{(}\sqrt{\delta S_{B_{x}}^{2}+\delta S_{B_{y}}^{2}}\Big{)}$ 0.54 nT/$\sqrt{\text{Hz}}$ 7.3 nT/$\sqrt{\text{Hz}}$
min$\big{(}S_{\theta}\big{)}$ 11 $\mu$rad/$\sqrt{\text{Hz}}$ 150 $\mu$rad/$\sqrt{\text{Hz}}$
8 MPE Drift Detection Without Recalibration
Running MPE calibrations described in Sec. 5 are used to compensate any microwave field drifts for the purposes of benchmarking vector magnetometry errors. In fact, all types of vector OPMs will inevitably experience drift in their reference system, whether defined by a coil system or an electromagnetic field, over extended periods. Current vector OPM techniques, however, have limited information from atomic measurements to detect sensor drift. For example, vector OPMs using coil modulations can only compare consistency between the norm of the measured vector components $(B_{x},B_{y},B_{z})$ with the scalar value $B$ to detect drift in parameters like coil non-orthogonality and coil factors. As a result, recalibration of the vector reference is often performed blindly, regardless of whether the drift occurred, causing additional downtime in the sensor acquisition.
In contrast to vector OPMs based on scalar detection, Rabi measurements across multiple transitions and MPEs contain much more independent information to detect drift in the MPE reference. For example, in this work there are nine independent Rabi measurements corresponding to the three spherical microwave components of MPE 1, MPE 2, and MPE 3. In comparison, if the frequency dependence of the microwave cavity modes is neglected, the sensor is characterized by 15 MPE parameters and two angles defining the DC magnetic field direction. However, because the microwave field corresponding to MPE 3 is the sum of the fields of MPE 1 and MPE 2, the number of independent MPE parameters can be further reduced. In this case, the number of parameters decreases from 15 to 11. This situation is expressed as
$\begin{split}\vec{\mathcal{B}}_{3}=&\vec{\mathcal{B}}_{1}+\vec{\mathcal{B}}_{2}=\{\mathcal{B}_{x,1}e^{-i\phi_{x,1}},\mathcal{B}_{y,1}e^{-i\phi_{y,1}},\mathcal{B}_{z,1}\}+\\ &\{\mathcal{B}_{x,2}e^{-i(\phi_{x,2}+\phi_{r})},\mathcal{B}_{y,2}e^{-i(\phi_{y,2}+\phi_{r})},\mathcal{B}_{z,2}e^{-i\phi_{r}}\},\end{split}$ (15)
where the additional parameter $\phi_{r}$ accounts for the relative phase between the microwave fields of MPE 1 and MPE 2. This is further demonstrated in Supplementary 1, where the independently calibrated MPE 3 parameters at a fixed microwave frequency are well approximated by the 10 microwave parameters of MPE 1 and MPE 2, along with a relative phase of $\phi_{r}=1.058$ rad, using Eq. (15). Thus, without frequency dependence of the cavity modes, the nine independent Rabi measurements are close to the 13 parameters (11 microwave and two angles defining the DC field orientation) that characterize the vector magnetometer. Two of these parameters will always be undefined without additional external information because any vector magnetometer cannot distinguish between global sensor rotations and DC magnetic field rotations.
Figure 6: The drift parameter $\delta_{\text{detect}}$ evaluated over 37.5 minutes demonstrating drift detection without full MPE recalibration.
As a proof of concept, Fig. 6 shows how the RMS error $\delta_{\text{detect}}=(r_{\text{rabi}}^{(\alpha,\beta)})^{1/2}$, computed from Eq. (9) after solving for $(\alpha,\beta)$ and setting the weights $(\tilde{w}_{m}^{m^{\prime}})_{s}\rightarrow 1$, provides a drift-detection measure. Establishing a specific threshold for $\delta_{\text{detect}}$ allows it to act as a cue to recalibrate, and thus eliminating the necessity for blind recalibration. Exploring how to further utilize MPE drift measures like $\delta_{\text{drift}}$ to mitigate MPE recalibration is the subject of future work.
9 Conclusion
This work demonstrates a vector OPM utilizing a microwave vector reference that achieves sub-milliradian accuracy [see Table 1] and a vector component sensitivity as low as 0.54 nT$/\sqrt{\text{Hz}}$ (11 $\mu$rad$/\sqrt{\text{Hz}}$) at magnetic field strengths near $50$ $\mu$T. This accuracy is reached by leveraging information in Rabi measurements across multiple hyperfine transitions to correct for systematic shifts arising from buffer gas collisions, nonlinear Zeeman effects, and off-resonant microwave driving. In addition, by using simultaneous precession and Rabi driving (SPaR), we also demonstrated deadzone-free vector measurements, even with the use of a single optical axis. Furthermore, utilizing three unique MPE configurations eliminated insensitive regions in the directional maps of the spherical microwave components [Fig. 1(c)] and enabled a unique measurement of the magnetic field direction, despite the periodicity of the spherical microwave components $\mathcal{B}_{k}^{(\alpha,\beta)}$ with respect to $(\alpha,\beta)$.
In future work, using magnetically quiet techniques like laser heating [62] to heat the vapor cell, rather than the entire microwave cavity, could eliminate measurement downtime and improve temperature stability, leading to more accurate MPE calibrations and longer durations of vector measurements before MPE drift errors become significant. Furthermore, designing a microwave system that does not depend on the microwave frequency, combined with full 3D control of the microwave field, could allow drift-tracking parameters derived from Rabi measurements (as discussed in Sec. 8) to detect and correct for microwave drifts, updating MPE parameters without the need for full recalibration via field rotations from the DC coil system.
The methodologies introduced here also lay the groundwork towards extending the Rabi vector measurements beyond its current scope. For instance, the techniques presented in Sec. 5 for calibrating the polarization ellipse of a microwave field could be useful for characterizing the accuracy of other microwave sensors based on atomic vapors such as atomic candle [63, 64, 65] and Rydberg EIT [66, 67, 68] methods. The Rabi vector technique could also be adapted with different electromagnetic references, such as radio-frequency fields that couple Zeeman transitions within a hyperfine manifold, commonly used in Mz and Mx configurations [69]. Alternatively, an all-optical approach could be explored by driving Rabi oscillations through two-photon Raman transitions. Furthermore, extending the model in Sec. 4 to also incorporate the amplitude of Rabi oscillations could be applied to the probing dead zone behavior [see Fig. 4] to accurately align the vector magnetometer frame to the orientation of the probe laser beam. This could have implications for referencing a vector OPM to non magnetic objects, vector gradiometry calibration, and tracking orientation drifts of the vector OPM not detectable with scalar calibration techniques [31].
[…]
Funding This work was supported by DARPA through ARO grant numbers W911NF-21-1-0127 and W911NF-19-1-0330, NSF QLCI Award OMA - 2016244, and the Baur-SPIE Chair in Optical Physics and Photonics at JILA.
[…]
Acknowledgments We acknowledge helpful conversations with Georg Bison, Michaela Ellmeier, Juniper Pollock, and Dawson Hewatt, and technical expertise from Yolanda Duerst, and Felix Vietmeyer.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
[…]
Supplemental document See Supplement 1 for supporting content.